OpenStudy (idealist10):

Find the function M such that M(x, y)dx+(x^2-y^2)dy=0 is exact.

OpenStudy (idealist10):

integral of (x^2-y^2)dy=yx^2-y^3/3+h(x)=C 2xy+h'(x)=M(x, y) But the answer is M(x, y)=2xy+f(x), how do I get to the answer?

OpenStudy (idealist10):

@jdoe0001 @goformit100 @nincompoop @e.mccormick

OpenStudy (idealist10):

@hartnn @amistre64

OpenStudy (idealist10):

I figured it out. Thanks for trying to help. :)